A project with a life of 8 years is expected to provide annual sales of $360,000 and costs of $253,000. The project will require an investment in equipment of $640,000, which will be depreciated on a straight-line method over the life of the project. You feel that both sales and costs are accurate to +/-15 percent. The tax rate is 40 percent. What is the annual operating cash flow for the best-case scenario?
Answer #1

Answer:

Best case scenario will be when:

Sales are higher by 15%

Costs are lower by 15%

Hence:

In best case scenario:

Annual Sales = 360000 * (1 + 15%) = $414,000

Annual Costs = 253000 * (1 - 15%) =  $215050

Annual depreciation =640000 / 8 = $80,000

Tax rate = 40%

Annual operating cash flow = (Sales - Costs) * (1 - Tax rate) + Depreciation * Tax rate

Hence:

Best case annual operating cash flow = (414000 - 215050) * (1 - 40%) + 80000 * 40% = $151370

Annual operating cash flow for the best-case scenario = $151,370